​(Common stock valuation​) Bates Inc. pays a dividend of ​$2.75 and is currently selling for ​$36.30. If investors require a return of 16 percent on their investment from buying Bates​ stock, what growth rate would Bates Inc. have to provide the​ investors?

The growth rate Bates Inc. would have to provide the investors is ??
Answer #1

Answer : Growth rate 7.83%

Required Return = (Expected Dividend / Current Price) + growth rate

0.16 = {2.75 * (1 + growth rate)] / 36.30} + growth rate

0.16 = 0.075757576 + 0.0757576 growth rate + growth rate

==>1.0757576 growth rate = 0.16 - 0.075757576

= 0.08424242425 / 1.0757575

= 0.0783 or 7.83%